in an ac current consisting of a resistor a capacitor and an inductor in series what happens to the resonance frequency if the resistance the capacitance the inductance and the amplitude are all double
Answer #1

Resonance\ frequency=\frac{1}{2\pi \sqrt{LC}}

Resonance frequency gets halved when inductance and capacitance is doubled. Resonance frequency does not depend on resistance and amplitude of the current.